OpenStudy (anonymous):

a population of 430 animals decreases at an annual rate of 21% find the multiplier for the rate of exponential decay

OpenStudy (amistre64):

i believe in general, (1+r) = k

OpenStudy (amistre64):

1 + r = .21 r = .21 - 1

OpenStudy (amistre64):

A = P(1+r)^x .21(430) = 430(1+r)^1 .21 = 1+r